What is the freezing point of 1.77 m solution of CBr4 in C6H6?
1
Expert's answer
2020-12-01T09:12:27-0500

We know that the equation is

ΔTf=Kf.Cm

Kf of benzene is =4.9k . molality-1

And Cm=1.77m

Hence ΔTf=4.9*1.77

=8.83k

And ΔTf=TF (pure solvent) − TF (solution)=8.83k

278.5k-TF(solution)=8.83k

TF(solution)=(278.5-8.83)k

=269.65k

Hence freezing point of solutions is 269.65k